OpenStudy (anonymous):

I GIVE MEDAL TO PEOPLE WHO HELPED ME Identify the property shown. (7 • p) • 2k = 7 • (p • 2k) A. distributive B. associative C. commutative D. transitive

OpenStudy (sleepyhead314):

distributive is when a ( b + c ) = a*b + a*c associative is when a + ( b + c ) = ( a + b ) + c [moving the parenthesis around] commutative is when a + b + c = a + c + b [switching the numbers into a different place] transitive is when if a = b and b = c , then a = c

OpenStudy (anonymous):

:D thx

OpenStudy (anonymous):

wait im still confused

OpenStudy (sleepyhead314):

what do you think your example is doing? (7 • p) • 2k = 7 • (p • 2k) hint: (a + b) + c = a + (b + c)

OpenStudy (sleepyhead314):

which option does it look like? :)

OpenStudy (anonymous):

is it A?

OpenStudy (sleepyhead314):

not quite :) that would have looked like (7 • p) • 2k = 14k * 2pk ^^^ which is not what happened ;P

OpenStudy (anonymous):

B ?

OpenStudy (sleepyhead314):

that's right! ^_^

OpenStudy (anonymous):

yay

OpenStudy (michele_laino):

hint: associative property with respect to the multiplication, is: \[\Large \left( {a \cdot b} \right) \cdot c = a \cdot \left( {b \cdot c} \right)\]

OpenStudy (anonymous):

thx u guys

OpenStudy (michele_laino):

:)

OpenStudy (sleepyhead314):

glad we could help :)

OpenStudy (anonymous):

:D